### v2

# Fourier restriction and well-approximable numbers

Abstract.
We use a deterministic construction to prove the optimality of the exponent in the Mockenhaupt-Mitsis-Bak-Seeger Fourier restriction theorem for dimension $d=1$ and parameter range $0<a,b\leq d$ and $b\leq 2a$. Previous constructions by Hambrook and Łaba [15] and Chen [5] required randomness and only covered the range $10<b\leq a\leq d=$ 1. We also resolve a question of Seeger [29] about the Fourier restriction inequality on the sets of well-approximable numbers.
1. Introduction
1.1. Statement of the problem
This paper concerns the optimality of the range of exponents $p$ in the following $L^{2}$ Fourier restriction theorem.
Theorem 1.1 (Mockenhaupt-Mitsis-Bak-Seeger).
Let $0<a,b<d$, with $d$ an integer. Define $p_{*}(a,b,d)=(4d-4a+2b)/b.$ Suppose $\mu$ is a Borel probability measure on $\mathbb{R}^{d}$ with the following properties:
  1. (A)
    $\mu(B(x,r))\lesssim r^{a}$ for all $x\in\mathbb{R}^{d}$ and all $r>0$.
  2. (B)
    $|\mathcal{F}(\mu)(\xi)|\lesssim(1+|\xi|)^{-b/2}$ for all $\xi\in\mathbb{R}^{d}$.
Then, for each $p\geq p_{*}(a,b,d)$, the following holds:
  1. (R)
    $\|\mathcal{F}(f\mu)\|_{L^{p}(\lambda)}\lesssim_{p}\|f\|_{L^{2}(\mu)}$ for all $f\in L^{2}(\mu)$.
Notation.
[…]
Mockenhaupt [25] and Mitsis [24] independently proved Theorem 1.1 with $p>p_{*}(a,b,d)$; the endpoint case $p=p_{*}(a,b,d)$ was established by Bak and Seeger [2]. Theorem 1.1 is a generalization of the classic Stein-Tomas Theorem [32], [33], which can be understood as the special case of Theorem 1.1 in which $\mu$ is the surface measure on a sphere in $\mathbb{R}^{d}$, $d\geq 2$. Generalizations of the Stein-Tomas Theorem to quadratic hypersurfaces and arbitrary smooth submanifolds in $\mathbb{R}^{d}$ were obtained, respectively, by Strichartz [31] and Greenleaf [13]; these results are also consequences of Theorem 1.1.
The following problem was suggested in the original papers of Mockenhaupt [25] and Mitsis [24].
Problem 1.2.
Given $a,b,d$, determine whether the value $p_{*}(a,b,d)$ in Theorem 1.1 is optimal.
To prove that $p_{*}(a,b,d)$ is optimal for a fixed triple $(a,b,d)$, we must prove the following: For each $p_{0}<p_{*}(a,b,d)$, there exists a Borel probability measure $\mu$ on $\mathbb{R}^{d}$ that satisfies (A) and (B) and there exists $p\geq p_{0}$ such that (R) is false.
Remark 1.3.
There are constructions of measures which satisfy (A), (B), and (R) for $p$ outside the range $p\geq p_{*}(a,b,d)$; see Shmerkin and Suomala [30], Chen and Seeger [6], Łaba and Wang [22], Ryou [27]. But this is a different problem than the one we consider here.
1.2. Motivation
As motivation for our main result, we start by discussing some counterexamples in the theory of restriction to curves and surfaces.
When $d\geq 2$ and $a=b=d-1$, $p_{*}(a,b,d)$ in Theorem 1.1 is known to be optimal because of a construction due (according to [33]) to Anthony W. Knapp. In Knapp’s construction, $\mu$ is the surface measure on a sphere, which satisfies (A) and (B) with $a=b=d-1$ (hence $p_{*}(a,b,d)=(2d+2)/(d-1)$). Knapp’s function $f$ is an indicator function of a sufficiently flat cap on the sphere. (To be more precise, $f$ is the indicator function of a rectangular region in $\mathbb{R}^{d}$ with center at a point on the sphere, with one side of length $\delta$, and with $d-1$ sides of length $\sqrt{\delta}$. The region is oriented so that the side of length $\delta$ lies normal to the sphere. The estimate (R) is shown to fail by taking $\delta\to 0$.)
In fact, Knapp’s example can be modified to apply for any smooth hypersurface. The following result appears as Lemma 3 of Strichartz [31] in the case $q=2$. It appears as Example 1.8 of Demeter [7] for general $q$.
Theorem 1.4.
Let $1\leq p,q\leq\infty.$ Let $S$ be a smooth $(d-1)$-dimensional submanifold of $\mathbb{R}^{d}$ and let $\mu$ be any smooth Borel probability measure on $S$. Then the estimate
$\|\mathcal{F}(f\mu)\|_{L^{p}(\lambda)}\lesssim_{p,q}\|f\|_{L^{q}(\mu)}\text{ for all $f\in L^{q}(\mu)$}$
cannot hold if $p<\frac{q(d+1)}{(q-1)(d-1)}.$
Note that if $a=b=d-1$ and $q=2$, this theorem recovers the exponent $p_{*}(a,b,d)$ from Theorem 1.1. Therefore, for the surface measure of a smooth hypersurface with non-vanishing Gaussian curvature, Theorem 1.1 is sharp.
Greenleaf [13, Remark 1.3] observed that Knapp’s example can be generalized to apply to any smooth hypersurface $S$ given as the Cartesian product of a $k$-dimensional surface and $(d-1-k)$-dimensional Euclidean space. In particular, if the $k$-dimensional surface has non-zero Gaussian curvature and if $\mu$ is the surface measure on $S$, this modified Knapp example shows that $p_{*}(a,b,d)$ in Theorem 1.1 is optimal when $a=d-1$ and $b=k$.
A measure $\mu$ satisfying (A) can only satisfy (R) if $p\geq 2d/a$. For a simple proof of this result, see Mitsis [24, Proposition 3.1]. This result does not tell us anything directly about Problem 1.2. However, by combining it with Theorem 1.1, it implies that if a measure satisfies (A) and (B), then $b\leq 2a$.
Ryou [27] builds on the $p\geq 2d/a$ result mentioned above using a Knapp-type example. In this way, Ryou establishes the following result for fractal subsets of the paraboloid (which appears as Theorem 1.3 in [27].)
Theorem 1.5.
Let $1\leq p,q\leq\infty.$ Let $\mathbb{P}^{d-1}$ denote the $(d-1)$-dimensional submanifold in $\mathbb{R}^{d}$ paramaterized by $(x,|x|^{2}):x\in\mathbb{R}^{d-1}$, and assume $\mu$ is a Borel probability measure supported on a subset of $\mathbb{P}^{d-1}$ of Hausdorff dimension $a\in(0,d-1)$ and satisfying the estimate
$\mu(B(x,r))\lesssim r^{a}$
for all $x\in\mathbb{P}^{d-1}$ and all $r>0$. Then the estimate
$\|\mathcal{F}(f\mu)\|_{L^{p}(\lambda)}\lesssim_{p,q}\|f\|_{L^{q}(\mu)}\text{ for all $f\in L^{q}(\mu)$}$
cannot hold if $p<\frac{q(d+1)}{(q-1)a}$.
Note that this result does not tell us about whether $p_{*}(a,b,d)$ is optimal in Theorem 1.1. However, by combining Theorem 1.1 and Theorem 1.5 when $q=2$, it follows that a measure $\mu$ that satisfies the hypothesis of Theorem 1.5 can only satisfy (B) with $b\leq 2a(d-a)/(1+d-a)$.
Restriction problems have also been considered for manifolds of codimension larger than $1$. A classic example is the moment curve $\Gamma\subseteq\mathbb{R}^{d}$ parameterized by $(x,x^{2},x^{3},\ldots,x^{d})$. The optimal extension estimates for this curve were obtained by Drury [8]:
Theorem 1.6.
Let $\mu$ denote the arc-length measure on the moment curve $\Gamma$. Then the estimate
$\|\mathcal{F}(f\mu)\|_{L^{p}(\lambda)}\lesssim_{p}\|f\|_{L^{q}(\mu)}\text{ forall $f\in L^{q}(\mu)$}$
holds if $p>\frac{d(d+1)}{2}+1$ and $q=\frac{2p}{2p-d(d+1)}$.
The endpoint $d(d+1)/2+1$ is known to be sharp; an argument of Arhipov, Cubarikov, and Karacuba [1] shows that Theorem 1.6 is false if $p=\frac{d(d+1)}{2}+1$. However, this argument does not tell us whether $p_{*}(a,b,d)$ is optimal in Theorem 1.1. Indeed, when we consider the arc-length measure on the moment curve, $a=1$, $b=1/d$, and $p_{*}(a,b,d)=4d(d-1)+2$, which is strictly larger than $d(d+1)/2+1$.
We now turn to past results on Problem 1.2 when $a$ and $b$ are not necessarily integers. Hambrook and Łaba [15] (see also Hambrook [14]) proved the optimality of $p_{*}(a,b,d)$ in Theorem 1.1 when $d=1$ and $0<b\leq a\leq d$. Since there is no notion of curvature in one dimension, the construction takes a form completely different from Knapp’s. The authors modified a construction of Łaba and Pramanik [23] to construct a random Cantor set containing a small deterministic Cantor set determined by a multi-scale arithmetic progression; more details can be found in Subsection 1.4 below. The construction was improved in technical ways by Chen [5]. In [17], Hambrook and Łaba proved the optimality of $p_{*}(a,b,d)$ for $d\geq 2$ and $d-1\leq b\leq a\leq d$. The construction combines the examples of Knapp and [15].
Seeger [29] asked whether it is possible to prove optimality of Theorem 1.1 when $d=1$ with a deterministic construction. In particular, Seeger wondered whether Theorem 1.1 is optimal for measures on the following sets:
$E(\alpha)=\left\{{x\in\mathbb{R}:|x-r/q|\leq|q|^{-(2+\alpha)}\text{ for infinitely many }(q,r)\in\mathbb{Z}\times\mathbb{Z}}\right\},$
where $\alpha>0$. The set $E(\alpha)$ is called the set of $\alpha$-well-approximable numbers. These sets arise from number theory (in particular, Diophantine approximation). They are important in harmonic analysis because they are the only known deterministic examples of so-called Salem sets in $\mathbb{R}$.
A set in $\mathbb{R}^{d}$ is called a Salem set if it supports a probability measure $\mu$ that satisfies (B) for every value of $b$ less than the Hausdorff dimension of the set. There are many random constructions of Salem sets in $\mathbb{R}^{d}$ (see [28], [20], [19], [18], [3], [6], [9], [23], [30]). Trivial examples of deterministic Salem sets in $\mathbb{R}^{d}$ are points (Hausdorff dimension 0), spheres (Hausdorff dimension $d-1$), and balls (Hausdorff dimension $d$). Kaufman [21] proved that $E(\alpha)$ is a Salem set of dimension $2/(2+\alpha)$ when $\alpha>0$. The sets $E(\alpha)$ are the only known deterministic examples of Salem sets in $\mathbb{R}$ of dimension other than 0 or 1.
1.3. Main Result
In the present paper, we prove that $p_{*}(a,b,d)$ is optimal for $d=1$, $0<a,b<d$ and $b\leq 2a$ with a completely deterministic construction. As we mentioned above, $b>2a$ cannot happen. Thus we cover all possible $a$ and $b$ when $d=1$. Recall that previous work only covered the case $b\leq a$. We also resolve the question of Seeger mentioned above. Our precise result is contained in the following two theorems.
Theorem 1.7.
Let $\alpha>0$ and $10\leq\beta<$ 1. There exists a Borel probability measure $\mu$ such that the following hold:
  • The support of $\mu$ is contained in $[-1/2,1/2]\cap E(\alpha)$.
  • For any $\epsilon>0$,
    $\mu(I)\lesssim_{\epsilon}|I|^{\frac{2}{2+\alpha}-\epsilon}\quad\text{for everyinterval $I\subseteq\mathbb{R}$}.$ (1.1)
  • For any $\epsilon>0$,
    $|\mathcal{F}(\mu)(\xi)|\lesssim_{\epsilon}(1+|\xi|)^{-\frac{1-\beta}{2+\alpha}+\epsilon}\quad\text{for every $\xi\in\mathbb{R}$}.$ (1.2)
  • There exists a sequence of nonnegative functions $\{f_{k}\}_{k\in\mathbb{N}}$ such that
    $\sup_{k\geq 1}\frac{{\left\|\mathcal{{F}}(f_{k}\mu)\right\|}_{L^{p}(\lambda)}}{{\left\|f_{k}\right\|}_{L^{q}(\mu)}}=\infty\qquad\text{whenever}\ p<p_{+}(q):=\frac{q}{q-1}\left(\frac{1+\alpha-\beta}{1-\beta}\right).$ (1.3)
Theorem 1.8.
Let $\alpha>0$ and $-1<\beta<0$. There exists a Borel probability measure $\mu$ such that the following hold:
  • The support of $\mu$ is contained in $[-1/2,1/2]\cap E(\alpha)$.
  • For any $\epsilon>0$,
    $\mu(I)\lesssim_{\epsilon}|I|^{\frac{2+\beta}{2+\alpha}-\epsilon}\quad\text{forevery interval $I\subseteq\mathbb{R}$}.$ (1.4)
  • For any $\epsilon>0$,
    $|\mathcal{F}(\mu)(\xi)|\lesssim_{\epsilon}(1+|\xi|)^{-\frac{1}{2+\alpha}+\epsilon}\quad\text{for every $\xi\in\mathbb{R}$}.$ (1.5)
  • There exists a sequence of nonnegative functions $\{f_{k}\}_{k\in\mathbb{N}}$ such that
    $\sup_{k\geq 1}\frac{{\left\|\mathcal{{F}}(f_{k}\mu)\right\|}_{L^{p}(\lambda)}}{{\left\|f_{k}\right\|}_{L^{q}(\mu)}}=\infty\qquad\text{whenever}\ p<p_{-}(q):=\frac{q}{q-1}\left(1+\alpha-\beta\right).$ (1.6)
Remark 1.9.
Theorem 1.7 and 1.8 imply the optimality of $p_{*}(a,b,d)$ in Theorem 1.1 when $10<a,b<d=$ 1 and $b\leq 2a$. This is a consequence of the following argument. Let $p_{0}<p_{*}(a,b,d)$ be given. Choose $\epsilon>0$ such that $p_{0}<p_{*}(a+\epsilon,b+\epsilon,d)<p_{*}(a,b,d)$.
If $b\leq a$, choose $\alpha>0$ and $10\leq\beta<$ 1 such that
$a+\epsilon=\frac{2}{2+\alpha},\qquad b+\epsilon=\frac{2(1-\beta)}{2+\alpha}.$
Set $q=2$. Then
$p_{+}(2)=2\left(\frac{1+\alpha-\beta}{1-\beta}\right)=\frac{4-4a+2b-2\epsilon}{b+\epsilon}=p_{*}(a+\epsilon,b+\epsilon,d).$
If $a<b\leq 2a$, choose $\alpha>0$ and $-1<\beta<0$ such that
$a+\epsilon=\frac{2+\beta}{2+\alpha},\qquad b+\epsilon=\frac{2}{2+\alpha}.$
Set $q=2$. Then
$p_{-}(2)=2(1+\alpha-\beta)=\frac{4-4a+2b-2\epsilon}{b+\epsilon}=p_{*}(a+\epsilon,b+\epsilon,d).$
[…]
Figure 1. If we ignore $\epsilon$ in the exponents, $(a,b)=(\frac{2}{2+\alpha},\frac{2}{2+\alpha})$ when $\beta=0$. The point $(a,b)\rightarrow(\frac{2}{2+\alpha},0)$ as $\beta\rightarrow 1$ and $(a,b)\rightarrow(\frac{1}{2+\alpha},\frac{2}{2+\alpha})$ as $\beta\rightarrow-1$.
1.4. Idea of the Proof of Theorem 1.7 and Theorem 1.8
The proof of Theorem 1.7 combines ideas from Kaufman [21], Papadimitropoulos [26], and Hambrook and Łaba [15]. We will first describe some key features of those works. Then will describe our proof of Theorem 1.7 and 1.8.
Kaufman [21] constructs a measure on $E(\alpha)$ that obeys (B) with $b=2/(2+\alpha)$ (here and elsewhere in this subsection, we ignore logarithmic factors and arbitrarily small exponent losses in (A) and (B)). However, as observed by Papadimitropoulos [26], Kaufman’s measure satisfies (A) with $a=1/(2+\alpha)$ but not any larger value of $a$. Papadimitropoulos [26] modifies Kaufman’s construction to obtain a measure that satisfies (A) and (B) with $a=b=2/(2+\alpha)$. Note that neither Kaufman nor Papadimitropoulos considered the inequality (R) for their measures. Note also that Kaufman’s and Papadimitropoulos’s measures are actually constructed on a Cantor set contained in $E(\alpha)$. As we describe below, this Cantor set plays an important role in our proof described below.
Hambrook and Łaba [15] start with a random Cantor set and then adjoin a deterministic Cantor set with arithmetic progression structure. More specifically, at each scale, they replace a small number of intervals of the random Cantor set with intervals of the same length but whose centers lie in a generalized arithmetic progression; these replacement intervals are the intervals of the structured deterministic Cantor set. The original random Cantor set is built so that its natural measure satisfies (A) and (B), for arbitrary prescribed values of $a$ and $b$ with $0<b\leq a\leq 1$. (In fact, Hambrook and Łaba [15] only obtained the case $a=b$. Chen [5] and Hambrook [14] modified the construction to get $b\leq a$.) Because the number of intervals replaced is small enough, the natural measure on the new Cantor set (i.e., the Cantor set obtained by adjoining the deterministic Cantor set to the random one) satisfies (A) and (B) with the same values of $a$ and $b$. The measure on the new Cantor set is the measure that shows $p_{*}(a,b,d)$ is optimal for Theorem 1.1. The functions $f_{k}$ which witness the failure of (R) are bump functions on the intervals in the $k$-th scale of the deterministic Cantor set. The arithmetic progression structure of the deterministic Cantor set is the key to obtaining the failure of (R).
For the proof of Theorem 1.7 and Theorem 1.8, we cannot simply add an arithmetic progression-structured Cantor set to $E(\alpha)$ because we need the support of the measure $\mu$ to be inside $E(\alpha)$. Instead, we must identify an appropriately structured set that is contained in $E(\alpha)$. We start with a Cantor set $\bigcap_{k\in\mathbb{N}}E(\alpha,k)$ contained in $E(\alpha)$; this is essentially the same Cantor set that Kaufman’s (and Papadimitropoulos’s) measure was constructed on. The precise definition of $E(\alpha,k)$ is given in Section 3. For now, it is enough to know that $E(\alpha,k)$ is a union of intervals of length $M_{k}^{-(2+\alpha)}$, where $M_{k}$ is a rapidly increasing sequence of positive numbers. At the $k$-th scale, if $k$ is odd, we adjoin to $E(\alpha,k)$ a set $C(\beta,k)$ which consists of a small number of intervals of length $M_{k}^{-(2+\alpha)}$ whose centers form an arithmetic progression. (The precise definition of $C(\beta,k)$ is given in Section 3.) For even $k$, we do not adjoin anything to $E(\alpha,k)$. We obtain a Cantor set
$\bigcap_{k\text{ even}}E(\alpha,k)\cap\bigcap_{k\text{ odd}}(E(\alpha,k)\cup C(\beta,k)).$
[…]
1.5. Structure of Remainder of Paper
The rest of the paper is devoted to the proof of Theorem 1.7 and Theorem 1.8. Section 2 states and proves the Convolution Stability Lemma, which is a fundamental tool in the verification of the various properties of the measure $\mu$. Section 3 constructs the measure $\mu$, proves that it has support contained in $[-1/2,1/2]\cap E(\alpha)$, and proves that it satisfies the required regularity and Fourier decay estimates on $\mu$ (namely, (1.1) and (1.2) if $10\leq\beta<$ 1 and (1.4) and (1.5) if $-1<\beta<0$). Section 4 constructs the sequence of functions $f_{k}$ and verifies the failure of the inequality (R) for that sequence by proving (1.3) if $10\leq\beta<$ 1 and (1.6) if $-1<\beta<0$.
1.6. Acknowledgements
We thank Andreas Seeger for posing his problem to us and hence inspiring this paper. We thank the anonymous referee for their helpful comments and suggestions which improved the presentation of this paper.
2. Convolution Stability Lemma
In this section, we state and prove the Convolution Stability Lemma (Lemma 2.1). Before that, we present some discussion of it.
In the course of the proof of Theorem 1.7 and 1.8, we will construct a sequence of functions $g_{k}$ with Fourier transforms that are large at well-separated scales. We will obtain the measure $\mu$ as a weak limit of measures $\mu_{k}$ with densities given by
$F_{0}(x)g_{1}(x)\ldots g_{k}(x),$
where $F_{0}$ is a suitable cutoff function to be described later. We will show that $\widehat{\mu}_{k-1}(s)$ is small unless $|s|$ lies below a certain threshold, and that $\widehat{g}_{k}(s)$ is equal to zero unless $|s|$ is larger than some other threshold. If these thresholds are sufficiently well-separated, there will be very little interference when we multiply $\mu_{k-1}(x)$ and $g_{k}(x)$ to arrive at $\mu_{k}(x)$. The Convolution Stability Lemma quantifies this.
We will use the Convolution Stability Lemma to prove the required regularity and Fourier decay estimates on $\mu$ and to prove the failure of the inequality (R).
Kaufman [21] uses a lemma (the only lemma in that paper) that is similar to the Convolution Stability Lemma to establish the Fourier decay estimate on the measure constructed in that paper. The works [3], [26], [34, Ch.9], [16], and [11] do likewise with similar lemmas. (Note that Papadimitropoulos [26] proves both a Fourier decay and regularity estimate on a measure. The Fourier decay estimate is established using a lemma similar to Kaufman’s, but the regularity estimate is proved using a counting argument.) The Convolution Stability Lemma in the present paper differs from these other lemmas in that it gives different estimates at different frequency scales. This furnishes a more precise understanding of behaviour of the measures $\mu_{k}$ at different frequency scales and was helpful in formulating our proofs. Note that [10] and [12] previously used versions of the Convolution Stability Lemma of the type used in the present paper (rather than the type used in Kaufman’s paper).
Lemma 2.1 (Convolution Stability Lemma).
Let $\alpha>0$. For $-1<\beta<1$, define
$N(\beta)=99\max(1,(1+\beta)^{-1}).$
Let $M_{k-1}$ and $M_{k}$ be real numbers satisfying
$\min\left(M_{k-1},M_{k-1}^{1+\beta}\right)\geq 100,$
$M_{k}\geq 4^{N(\beta)+1}M_{k-1}^{(2+\alpha)(N(\beta)+1)}(1+M_{k-1}^{\beta}),$ (2.1)
$\log(M_{k})\geq N(\beta)M_{k-1}^{3(2+\alpha)}.$ (2.2)
Let $G,H:\mathbb{Z}\to\mathbb{C}$ be such that $|G|\leq 1$, $G(0)=1$, and $|H|\leq 2$. Assume $G$ and $H$ satisfy the following estimates:
$G(s)=0\qquad\text{if}\ 1\leq|s|<\min(M_{k},M_{k}^{1+\beta});$ (G1)
there is a constant $C>0$ such that
$|G(s)|\leq C\log(|s|)M_{k}^{-1}(1+M_{k}^{\beta})\qquad\text{if}\ |s|\geq\min(M_{k},M_{k}^{1+\beta});$ (G2)
for each integer $N\geq 1$ there is a constant $C_{1,N}>0$ such that
$|G(s)|\leq C_{1,N}\log(|s|)M_{k}^{-1+(2+\alpha)N}(1+M_{k}^{\beta})|s|^{-N}\qquad\text{if}\ |s|\geq M_{k}^{2+\alpha};$ (G3)
for each integer $N\geq 1$ there is a constant $C_{2,N}>0$ such that
$|H(s)|\leq C_{2,N}\log(|s|)M_{k-1}^{-1+(2+\alpha)(N+1)}(1+M_{k-1}^{\beta})|s|^{-N}\qquad\text{if}\ |s|\geq 2M_{k-1}^{2+\alpha}.$ (H)
Then the following estimates hold:
$|G\ast H(s)-H(s)|<C_{2,N(\beta)}M_{k}^{-96}\qquad\text{if}\ |s|\leq\frac{1}{2}\min(M_{k},M_{k}^{1+\beta});$ (GH1)
$|G\ast H(s)|\leq(C+1)(C_{2,N(\beta)}+1)\log^{2}(|s|)M_{k}^{-1}(1+M_{k}^{\beta})\qquad\text{if}\ |s|\geq\frac{1}{2}\min(M_{k},M_{k}^{1+\beta});$ (GH2)
$|G\ast H(s)|\leq(4^{N+1}C_{1,N+1}+C_{2,N+1})\log(|s|)M_{k}^{-1+(2+\alpha)(N+1)}(1+M_{k}^{\beta})|s|^{-N}$ (GH3)
$\hskip 260.17464pt\text{if}\ |s|\geq 2M_{k}^{2+\alpha}.$
Remark 2.2.
The indices on $M_{k-1}$ and $M_{k}$ in Lemma 2.1 have no meaning at this point. We have used the notation because Lemma 2.1 will be used iteratively later. In particular, note that the estimate (GH3) is of the same form as the estimate (H) with $k-1$ replaced by $k$.
We will prove the estimates on $G\ast H$ separately.
Proof of (GH1).
Assume $|s|\leq\min(M_{k},M_{k}^{1+\beta})/2$. We use (G1), that $G(0)=1$, and that $|G|\leq 1$ to obtain
$|G\ast H(s)-H(s)|=\sum_{|t|\geq\min(M_{k},M_{k}^{1+\beta})}|G(t)H(s-t)|\leq\sum_{|t|\geq\min(M_{k},M_{k}^{1+\beta})}|H(s-t)|.$
Since $|t|\geq\min(M_{k},M_{k}^{1+\beta})$ and $|s|\leq\min(M_{k},M_{k}^{1+\beta})/2$, $|s-t|\geq|t|-|s|\geq\min(M_{k},M_{k}^{1+\beta})/2$. By (2.1), $\min(M_{k},M_{k}^{1+\beta})/2\geq 2M_{k-1}^{2+\alpha}$, and we can use (H) with $N=N(\beta)$. Therefore,
$\begin{split}|G\ast H(s)-H(s)|&\leq\sum_{|u|\geq\min(M_{k},M_{k}^{1+\beta})/2}|H(u)|\\ &\leq C_{2,N(\beta)}M_{k-1}^{-1+(2+\alpha)(N(\beta)+1)}(1+M_{k-1}^{\beta})\sum_{|u|\geq\min(M_{k},M_{k}^{1+\beta})/2}\log(|u|)|u|^{-N(\beta)}\\ &\leq 2\cdot C_{2,N(\beta)}\log(M_{k})M_{k-1}^{-1+(2+\alpha)(N(\beta)+1)}(1+M_{k-1}^{\beta})4^{N(\beta)}M_{k}^{-98}\\ &\leq C_{2,N(\beta)}M_{k}^{-96}.\end{split}$
In the last inequality, we used (2.1) and that $\log(|s|)\leq\log(M_{k})\leq M_{k}$. ∎
Proof of (GH2).
Assume $|s|\geq\min(M_{k},M_{k}^{1+\beta})/2$. We write
$|G\ast H|=\sum_{|t|<2M_{k-1}^{2+\alpha}}|G(s-t)H(t)|+\sum_{|t|\geq 2M_{k-1}^{2+\alpha}}|G(s-t)H(t)|.$
When $|t|<2M_{k-1}^{2+\alpha}$, the assumptions $|s|\geq\min(M_{k},M_{k}^{1+\beta})/2$ and (2.1) imply that $s-t\neq 0$ and $|s-t|\leq 2|s|$. Using (G1), (G2) and the assumption that $|H|\leq 2$, we get
$\sum_{|t|<2M_{k-1}^{2+\alpha}}|G(s-t)H(t)|\leq 2C\log(2|s|)M_{k}^{-1}(1+M_{k}^{\beta})(4M_{k-1}^{2+\alpha}+1).$
By (2.2),
$\sum_{|t|<2M_{k-1}^{2+\alpha}}|G(s-t)H(t)|\leq C\log^{2}(|s|)M_{k}^{-1}(1+M_{k}^{\beta}).$ (2.3)
When $|t|\geq 2M_{k-1}^{2+\alpha}$, we consider the cases $t=s$ and $t\neq s$ separately. If $t=s$, we use $G(0)=1$ and (H) with $N=N(\beta)$ to get
$\begin{split}|G(0)H(s)|&\leq C_{2,N(\beta)}\log(|s|)M_{k-1}^{-1+(2+\alpha)(N(\beta)+1)}(1+M_{k-1}^{\beta})|s|^{-N(\beta)}\\ &\leq C_{2,N(\beta)}2^{N(\beta)}\log(|s|)M_{k-1}^{-1+(2+\alpha)(N(\beta)+1)}(1+M_{k-1}^{\beta})M_{k}^{-99}.\end{split}$
Using (2.1), we obtain
$|G(0)H(s)|\leq C_{2,N(\beta)}\log(|s|)M_{k}^{-98}.$ (2.4)
If $t\neq s$, we use (G1), (G2), and (H) with $N=N(\beta)$ to get
$\begin{split}\sum_{\begin{subarray}{c}|t|\geq 2M_{k-1}^{2+\alpha}\\ t\neq s\end{subarray}}|G(s-t)H(t)|\leq&C\cdot C_{2,N(\beta)}M_{k}^{-1}(1+M_{k}^{\beta})M_{k-1}^{-1+(2+\alpha)(N(\beta)+1)}(1+M_{k-1}^{\beta})\\ &\times\sum_{\begin{subarray}{c}|t|\geq 2M_{k-1}^{2+\alpha}\\ t\neq s\end{subarray}}\log(|s-t|)\log(|t|)|t|^{-N(\beta)}.\end{split}$
Since $|t|\geq 2M_{k-1}^{2+\alpha}$, $\log^{2}(|t|)\leq|t|$. Thus, we have
$\begin{split}\sum_{\begin{subarray}{c}|t|\geq 2M_{k-1}^{2+\alpha}\\ t\neq s\end{subarray}}\log(|s-t|)\log(|t|)|t|^{-N(\beta)}&\leq\log(|s|)\sum_{|t|\geq 2M_{k-1}^{2+\alpha}}\log^{2}(|t|)|t|^{-N(\beta)}\\ &\leq\log(|s|)\sum_{|t|\geq 2M_{k-1}^{2+\alpha}}|t|^{-N(\beta)+1}\\ &\leq\log(|s|)M_{k-1}^{-(2+\alpha)(N(\beta)-2)}.\end{split}$
Therefore,
$\sum_{\begin{subarray}{c}|t|\geq 2M_{k-1}^{2+\alpha}\\ t\neq s\end{subarray}}|G(s-t)H(t)|\leq C\cdot C_{2,N(\beta)}\log(|s|)M_{k}^{-1}(1+M_{k}^{\beta})M_{k-1}^{-1+3(2+\alpha)}(1+M_{k-1}^{\beta}).$
Since $|s|\geq\min(M_{k},M_{k}^{1+\beta})/2$, by (2.2), we obtain
$\sum_{\begin{subarray}{c}|t|\geq 2M_{k}^{2+\alpha}\\ t\neq s\end{subarray}}|G(s-t)H(t)|\leq C\cdot C_{2,N(\beta)}\log^{2}(|s|)M_{k}^{-1}(1+M_{k}^{\beta}).$ (2.5)
Combining (2.3), (2.4) and (2.5), we get (GH2). ∎
Proof of (GH3).
We write
$\begin{split}|G\ast H(s)|=\sum_{|t|\leq|s|/2}|G(t)H(s-t)|+\sum_{|t|\geq|s|/2}|G(t)H(s-t)|.\end{split}$
When $|t|\leq|s|/2$, we have $|t-s|\geq|s|-|t|\geq|s|/2$. Since $|G|\leq 1$,
$\sum_{|t|\leq|s|/2}|G(t)H(s-t)|\leq\sum_{|t|\leq|s|/2}|H(s-t)|\leq\sum_{|u|\geq|s|/2}|H(u)|.$
Since $|s|\geq 2M_{k}^{2+\alpha}$, the assumption (2.1) implies that $|s|/2\geq 2M_{k-1}^{2+\alpha}$, so we can use (H). By using (H) (with $N+1$ in place of $N$), we obtain
$\begin{split}\sum_{|t|\leq|s|/2}|G(t)H(s-t)|&\leq C_{2,N+1}M_{k-1}^{-1+(2+\alpha)(N+2)}(1+M_{k-1}^{\beta})\sum_{|u|\geq|s|/2}\log(|u|)|u|^{-(N+1)}\\ &\leq C_{2,N+1}\log(|s|)M_{k-1}^{-1+(2+\alpha)(N+2)}(1+M_{k-1}^{\beta})2\cdot 4^{N}|s|^{-N}.\end{split}$
By using (2.1), we get
$\begin{split}2\cdot 4^{N}M_{k-1}^{-1+(2+\alpha)(N+2)}(1+M_{k-1}^{\beta})&\leq 4^{N+1}M_{k-1}^{(2+\alpha)(N+2)}\\ &\leq M_{k}^{(N+2)/100}\\ &\leq M_{k}^{-1+(2+\alpha)(N+1)}(1+M_{k}^{\beta}).\end{split}$
Thus,
$\sum_{|t|\leq|s|/2}|G(t)H(s-t)|\leq C_{2,N+1}\log(|s|)M_{k}^{-1+(2+\alpha)(N+1)}(1+M_{k}^{\beta})|s|^{-N}.$
When $|t|\geq|s|/2\geq M_{k}^{2+\alpha}$, we use (G3) (with $N+1$ in place of $N$) and that $|H|\leq 2$. Thus
$\begin{split}\sum_{|t|\geq|s|/2}|G(t)H(s-t)|&\leq 2\sum_{|t|\geq|s|/2}|G(t)|\\ &\leq 2C_{1,N+1}M_{k}^{-1+(2+\alpha)(N+1)}(1+M_{k}^{\beta})\sum_{|t|\geq|s|/2}\log(|t|)|t|^{-(N+1)}\\ &\leq 2C_{1,N+1}\log(|s|)M_{k}^{-1+(2+\alpha)(N+1)}(1+M_{k}^{\beta})2\cdot 4^{N}|s|^{-N}.\end{split}$
Combining the two cases, we get (GH3). ∎
3. Construction and properties of the measure $\mu$
Let $\alpha>0$ and $-1<\beta<1$ be given. In this section, we will construct the measure $\mu$, prove that it has support contained in $\mathbb{T}\cap E(\alpha)$, and prove that it satisfies
  • (1.1) and (1.2) in case $10\leq\beta<$ 1, or
  • (1.4) and (1.5) in case $-1<\beta<0$.
3.1. Preliminaries
[…]
Since $F$ and $F_{0}$ are Schwartz functions, there are nondecreasing sequences of numbers $C_{N},\widetilde{C}_{N}\geq 1$ (where $N=0,1,2\ldots$) such that
$|\widehat{F}(s)|\leq C_{N}|s|^{-N}\quad\text{ and }\quad|\widehat{F}_{0}(s)|\leq\widetilde{C}_{N}|s|^{-N}$ (3.1)
for each $|s|\geq 1$. Indeed, one may take
$C_{N}=\max_{0\leq i\leq N}\sup_{x\in\mathbb{T}}|\partial^{i}F|\qquad\text{and}\qquad\widetilde{C}_{N}=\max_{0\leq i\leq N}\sup_{x\in\mathbb{T}}|\partial^{i}F_{0}|.$
Let $\alpha>0$ and $-1<\beta<1$. We fix a sequence $\{\epsilon_{k}\}_{k\in\mathbb{N}}$ that converges to 00. For each $k$, we choose $N_{k}$ such that $N_{k}\geq\max(N(\beta),k)$ and $N_{k}\epsilon_{k}\geq 2+\alpha$.
Let $P_{M}$ denote the set of prime numbers strictly between $M$ and $2M$. Fix $D>1$. We choose a rapidly increasing sequence of numbers $\{M_{k}\}_{k\in\mathbb{N}\cup\{0\}}$ that satisfies the following:
  • $\min(M_{0},M_{0}^{1+\beta},M_{0}^{\alpha})\geq 100$ and $\min(M_{0},M_{0}^{1-\beta})\geq\log(M_{0})$.
  • For any $k$,
    $\frac{M_{k}}{{D}\log(M_{k})}\leq|P_{M_{k}}|.$
  • For any $k$,
    $\log(M_{k})\geq 4^{N(\beta)+1}N(\beta)M_{k-1}^{(2+\alpha)(N(\beta)+1)}.$ (3.2)
  • For any $k$,
    $\log\left(\frac{1}{2}\min(M_{k},M_{k}^{1+\beta})\right)\geq\left(\frac{1}{1+\beta}+1\right)^{2}(1+2{D})\left(1+\widetilde{C}_{N_{k}+k}+8^{N_{k}+k+2}C_{N_{k}+k}D\ \right).$ (3.3)
Note that (3.2) implies (2.1) and (2.2) in Lemma 2.1 since $M_{k}\geq\log(M_{k})$.
Recall that
$E(\alpha)=\left\{{x\in\mathbb{R}:|x-r/q|\leq|q|^{-(2+\alpha)}\text{ for infinitely many }(q,r)\in\mathbb{Z}\times\mathbb{Z}}\right\}.$
For each $k\in\mathbb{N}$, define
$E(\alpha,k)=\left\{{x\in\mathbb{R}:|x-r/p|\leq\frac{M_{k}^{-(2+\alpha)}}{2}\text{ for some }p\in P_{M_{k}},r\in\mathbb{Z},|r|\leq(p-1)/2}\right\}.$
Note that $E(\alpha,k)$ is a subset of $\mathbb{T}$. Note also that any point belonging to $E(\alpha,k)$ for infinitely many $k$ will belong to $E(\alpha)$.
For each odd $k\in\mathbb{N}$, we choose a number $B_{k}\in\mathbb{N}$ such that $M_{k}^{1+\beta}\leq B_{k}\leq 2M_{k}^{1+\beta}$. Additionally, if $\beta=0$, we choose $B_{k}\in P_{M_{k}}$; if $\beta\neq 0$, we choose $B_{k}$ such that $B_{k}$ is coprime to every $p\in P_{M_{k}}$ for all sufficiently large $k$.
For each odd $k\in\mathbb{N}$, define
$C(\beta,k)=\left\{{x\in\mathbb{R}:\left|x-\frac{a}{B_{k}}\right|\leq\frac{M_{k}^{-(2+\alpha)}}{2}\ \text{ for some }a\in\mathbb{Z},|a|\leq(B_{k}-1)/2}\right\}.$
Note that $C(\beta,k)$ is a subset of $\mathbb{T}$.
Remark 3.1.
For $\beta=0$, it is possible to make the same choice of $B_{k}$ as we do in the $\beta\neq 0$ case. However, the proof is somewhat simpler and more intuitive in the $\beta=0$ case with the choice we have made.
3.2. The single-scale functions $g_{k}$
For even $k$, define
$g_{k}(x):=\frac{M_{k}^{1+\alpha}}{|P_{M_{k}}|}\sum_{p\in P_{M_{k}}}\sum_{|r|\leq(p-1)/2}F(M_{k}^{1+\alpha}(px-r)).$
For odd $k$, define
$\begin{split}g_{k}(x):=&\frac{M_{k}^{1+\alpha}}{|P_{M_{k}}|+M_{k}^{\beta}+1}\sum_{p\in P_{M_{k}}}\sum_{|r|\leq(p-1)/2}F(M_{k}^{1+\alpha}(px-r))\\ &+\frac{M_{k}^{1+\alpha}(1+M_{k}^{-\beta})}{|P_{M_{k}}|+M_{k}^{\beta}+1}\sum_{|a|\leq(B_{k}-1)/2}F(M_{k}^{1+\alpha-\beta}(B_{k}x-a)).\end{split}$
[…]
Lemma 3.2.
For every $k\in\mathbb{N}$ and $s\in\mathbb{Z}$, the following statements hold.
  1. (1)
    $|\widehat{g}_{k}(s)|\leq 1$.
  2. (2)
    $\widehat{g}_{k}(s)=0$ if $1\leq|s|<\min(M_{k},M_{k}^{1+\beta})$.
  3. (3)
    If $s\neq 0$,
    $|\widehat{g}_{k}(s)|\leq 2D\left(\frac{1}{1+\beta}+1\right)\log(|s|)M_{k}^{-1}(1+M_{k}^{\beta}).$ (3.4)
  4. (4)
    If $s\neq 0$,
    $|\widehat{g}_{k}(s)|\leq 2^{N+1}C_{N}D\left(\frac{1}{1+\beta}+1\right)\log(|s|)M_{k}^{-1+(\alpha+2)N}(1+M_{k}^{\beta})|s|^{-N}.$ (3.5)
Proof.
Since $F$ and $g_{k}$ are supported on $\mathbb{T}$, $\widehat{F}(s)=\mathcal{F}(F)(s)$ and $\widehat{g_{k}}(s)=\mathcal{F}(g_{k})(s)$. We shall need the following calculation. Suppose $c>0$ and $b$ is an odd positive integer. In the following, the variable $a$ is an integer. For every $s\in\mathbb{Z}$,
$\int_{\mathbb{R}}\sum_{|a|\leq(b-1)/2}F(c(bx-a))e^{-2\pi ixs}dx=\dfrac{1}{cb}\widehat{F}\left({\dfrac{s}{cb}}\right)e^{\pi i(b-1)s/b}\sum_{a=0}^{b-1}e^{-2\pi ias/b}$
The sum on the right equals $b$ if $b$ divides $s$ and equals 00 otherwise. We now apply this calculation to compute $\widehat{g}_{k}$ and prove (1), (2), (3), and (4).
First assume $k$ is even. We have
$\widehat{g}_{k}(s)=\frac{1}{|P_{M_{k}}|}\sum_{p\in P_{M_{K}}}e^{\pi i(p-1)s/p}\widehat{F}\left(\frac{s}{pM_{k}^{1+\alpha}}\right)\mathbf{1}_{p\mathbb{Z}}(s),$
where $\mathbf{1}_{p\mathbb{Z}}(s)=1$ if $p\mid s$ and 00 otherwise. Hence (1) and (2) follow easily. It remains to establish (3) and (4). By (3.1), we have the estimate
$\left|\widehat{F}\left(\frac{s}{pM_{k}^{1+\alpha}}\right)\right|\leq\min(1,2^{N}C_{N}M_{k}^{(2+\alpha)N}|s|^{-N}).$
Since we have already established (2) for $1\leq|s|<M_{k}$, it is enough to consider $|s|\geq M_{k}$. Since there are at most $\frac{\log|s|}{\log M_{k}}$ values of $p\in P_{M_{k}}$ such that $p\mid s$, we have
$|\hat{g}_{k}(s)|$ $\leq DM_{k}^{-1}\bigg{(}\log(|s|)+\log(M_{k})\bigg{)}\min(1,2^{N}C_{N}M_{k}^{(2+\alpha)N}|s|^{-N})$
$\leq 2DM_{k}^{-1}\log(|s|)\min(1,2^{N}C_{N}M_{k}^{N(2+\alpha)}|s|^{-N}),$
which immediately gives (3) and (4).
Now consider the case where $k$ is odd. We have
$\begin{split}\widehat{g}_{k}(s)=&\frac{1}{|P_{M_{k}}|+M_{k}^{\beta}+1}\sum_{p\in P_{M_{k}}}e^{\pi i(p-1)s/p}\widehat{F}\left(\frac{s}{pM_{k}^{1+\alpha}}\right)\mathbf{1}_{p\mathbb{Z}}(s)\\ &+\frac{e^{\pi i(B_{k}-1)s/B_{k}}(1+M_{k}^{\beta})}{|P_{M_{k}}|+M_{k}^{\beta}+1}e^{\pi i(B_{k}-1)/B_{k}}\widehat{F}\left(\frac{s}{B_{k}M_{k}^{\alpha-\beta+1}}\right)\mathbf{1}_{B_{k}\mathbb{Z}}(s),\end{split}$ (3.6)
where $\mathbf{1}_{p\mathbb{Z}}(s)$ is defined as before and $\mathbf{1}_{B_{k}\mathbb{Z}}(s)=1$ if $B_{k}|s$ and $\mathbf{1}_{B_{k}\mathbb{Z}}(s)=0$ otherwise. Thus, (1) and (2) easily follow. By (3.1),
$\left|{\widehat{F}\left(\frac{s}{pM_{k}^{1+\alpha}}\right)}\right|\leq\min(1,2^{N}C_{N}M_{k}^{(2+\alpha)N}|s|^{-N})$
and
$\left|{\widehat{F}\left(\frac{s}{B_{k}M_{k}^{\alpha-\beta+1}}\right)}\right|\leq\min(1,2^{N}C_{N}M_{k}^{(2+\alpha)N}|s|^{-N}).$
If $\beta\geq 0$, $\widehat{g}_{k}(s)=0$ if $1\leq s<M_{k}$ by (2). Thus, it suffices to consider $|s|\geq M_{k}$. Using that the number of $p\in P_{M_{k}}$ such that $p|s$ is bounded by $\frac{\log|s|}{\log M_{k}}$, we obtain
$\begin{split}|\widehat{g}_{k}(s)|&\leq DM_{k}^{-1}\left(\log(|s|)+\log(M_{k})(1+M_{k}^{\beta})\right)\min(1,2^{N}C_{N}M_{k}^{(2+\alpha)N}|s|^{-N})\\ &\leq 2DM_{k}^{-1}\log(|s|)(1+M_{k}^{\beta})\min(1,2^{N}C_{N}M_{k}^{N(2+\alpha)}|s|^{-N}).\end{split}$
If $-1<\beta<0$, it suffices to consider $|s|\geq M_{k}^{1+\beta}$. Similarly, we obtain that
$|\widehat{g}_{k}(s)|\leq DM_{k}^{-1}\log(|s|)\left(1+\frac{1}{1+\beta}(1+M_{k}^{\beta})\right)\min(1,2^{N}C_{N}M_{k}^{(2+\alpha)N}|s|^{-N}).$
Therefore, we get (3) and (4). ∎
3.3. The measures $\mu_{k}$
We define a sequence of functions
$\mu_{k}=g_{k}\cdots g_{1}F_{0}$
and we identify $\mu_{k}$ with the measure $\mu_{k}dx$.
Lemma 3.3.
The following statements hold for all $k\in\mathbb{N}$ and $s\in\mathbb{Z}$:
  1. (1)
    We have
    $|\widehat{\mu}_{k}(s)|\leq 2.$ (3.7)
  2. (2)
    If $|s|\leq\min(M_{k},M_{k}^{1+\beta})/2$,
    $|\widehat{\mu}_{k}(s)-\widehat{\mu}_{k-1}(s)|\leq\log(M_{k})M_{k}^{-96}.$ (3.8)
  3. (3)
    If $|s|\geq\min(M_{k},M_{k}^{1+\beta})/2$,
    $|\widehat{\mu}_{k}(s)|\leq\log^{3}(|s|)M_{k}^{-1}(1+M_{k}^{\beta}).$ (3.9)
  4. (4)
    If $|s|\geq 2M_{k}^{2+\alpha}$, for any $N\in\mathbb{N}$,
    $|\widehat{\mu}_{k}(s)|\leq\left(\widetilde{C}_{N+k}+2D\left(\frac{1}{1+\beta}+1\right)\sum_{j=1}^{k}8^{N+j}C_{N+j}\right)\log(|s|)M_{k}^{-1+(2+\alpha)(N+1)}(1+M_{k}^{\beta})|s|^{-N}.$ (3.10)
Proof.
First, we prove by induction that the following hold for all $k\in\mathbb{N}$:
  1. (i)
    If $|s|\leq\min(M_{k},M_{k}^{1+\beta})/2$,
    $|\widehat{\mu}_{k-1}\ast\widehat{g}_{k}(s)-\widehat{\mu}_{k-1}(s)|\leq\left(\widetilde{C}_{N(\beta)+k-1}+2D\left(\frac{1}{1+\beta}+1\right)\sum_{j=1}^{k}8^{N(\beta)+j-1}C_{N(\beta)+j-1}\right)M_{k}^{-96}.$
  2. (ii)
    If $|s|\geq\min(M_{k},M_{k}^{1+\beta})/2$,
    $\begin{split}|\widehat{\mu}_{k-1}\ast\widehat{g}_{k}(s)|\leq&\left(\frac{1}{1+\beta}+1\right)^{2}\left(1+2{D}\right)\left(1+\widetilde{C}_{N(\beta)+k-1}+2D\sum_{j=1}^{k}8^{N(\beta)+j-1}C_{N(\beta)+j-1}\right)\\ &\times\log^{2}(|s|)M_{k}^{-1}(1+M_{k}^{\beta}).\end{split}$
  3. (iii)
    If $|s|\geq 2M_{k}^{2+\alpha}$, for any $N\in\mathbb{N}$,
    $\begin{split}|\widehat{\mu}_{k-1}\ast\widehat{g}_{k}(s)|\leq&\left(\widetilde{C}_{N+k}+2D\left(\frac{1}{1+\beta}+1\right)\sum_{j=1}^{k}8^{N+j}C_{N+j}\right)\\ &\times\log(|s|)M_{k}^{-1+(2+\alpha)(N+1)}(1+M_{k}^{\beta})|s|^{-N}.\end{split}$
When $k=1$, $\widehat{\mu}_{0}=\widehat{F}_{0}$ where $F_{0}$ is a Schwartz function. Let $G=\widehat{g}_{1}$ with $C=2D((1+\beta)^{-1}+1)$ and $C_{1,N}=2^{N+1}C_{N}D((1+\beta)^{-1}+1)$ and let $H=\widehat{\mu}_{0}$ with $C_{2,N}=\widetilde{C}_{N}$. By Lemma 3.2, $\widehat{g}_{1}$ satisfies the assumptions on $G$ in Lemma 2.1. Because $|\widehat{\mu}_{0}(s)|\leq 1$ and because of (3.1), $\widehat{\mu}_{0}$ satisfies the assumptions on $H$ in Lemma 2.1. Thus, we have (i), (ii) and (iii) when $k=1$.
Now we assume (i), (ii), (iii) when $k=1,2,\cdots,i-1$ and prove the same statements with $i-1$ replaced by $i$. Let $G=\widehat{g}_{i}$ with $C=2D((1+\beta)^{-1}+1)$ and $C_{1,N}=2^{N+1}C_{N}D((1+\beta)^{-1}+1)$ as before, and let $H=\widehat{\mu}_{i-1}$ with $C_{2,N}=\widetilde{C}_{N+i-1}+2D((1+\beta)^{-1}+1)\sum_{j=1}^{i-1}8^{N+j}C_{N+j}$. By Lemma 3.2, $\widehat{g}_{i}$ satisfies the assumptions on $G$ in Lemma 2.1.
Note that $|\widehat{\mu}_{i-1}(s)|\leq|\widehat{\mu}_{i-1}(0)|$ since $\mu_{i-1}$ is nonnegative. The condition (3.2) implies that $M_{k}\geq\log(M_{k})\geq 2^{k}$ and (3.3) implies that
$\begin{split}\left(\widetilde{C}_{N(\beta)+k}+{2D}\left(\frac{1}{1+\beta}+1\right)\sum_{j=1}^{k}8^{N(\beta)+j}C_{N(\beta)+j}\right)&\leq\log\left(\frac{1}{2}\min(M_{k},M_{k}^{1+\beta})\right)\\ &\leq\log(M_{k}).\end{split}$ (3.11)
Thus, we use (i) for $k=1,2,\cdots,i-1$ and obtain
$|\widehat{\mu}_{i-1}(0)|\leq|\widehat{\mu}_{0}(0)|+\sum_{j=1}^{i-1}\log(M_{i})M_{i}^{-96}\leq 2.$ (3.12)
Also, applying (iii) with $k=i-1$ shows that $\hat{\mu}_{i-1}(s)$ satisfies the assumption (H) with $k=i$. Therefore, $\widehat{\mu}_{i-1}$ satisfies the assumptions on $H$ in Lemma 2.1. Thus, (i), (ii) and (iii) with $k=i$ hold by Lemma 2.1. By induction, (i),(ii) and (iii) hold for any $k\in\mathbb{N}$.
We already proved (4) by (iii). Combining (i) and (3.11), we get (2). Similarly, (3) follows from (ii) and (3.3). Since (3.12) holds for any $i\in\mathbb{N}$, we get (1). ∎
3.4. Definition and support of $\mu$
[…]
From the definition of $\mu_{k}$, it is clear the support of $\mu$ is contained in $\mathbb{T}\cap\bigcap_{k\in\mathbb{N}}\operatorname*{supp}(g_{k})$. But recall that the support of $g_{k}$ is contained in $E(\alpha,k)$ if $k$ is even and is contained in $E(\alpha,k)\cup C(\beta,k)$ if $k$ is odd. Recall also that any point that belongs to $E(\alpha,k)$ for infinitely many $k$ must belong to $E(\alpha)$. Therefore
$\bigcap_{k\in\mathbb{N}}\operatorname*{supp}(g_{k})\subseteq\bigcap_{k\text{ even}}E(\alpha,k)\cap\bigcap_{k\text{ odd}}\left({E(\alpha,k)\cup C(\beta,k)}\right)\subseteq\bigcap_{k\text{ even}}E(\alpha,k)\subseteq E(\alpha).$
This shows that $\operatorname*{supp}(\mu)\subseteq\mathbb{T}\cap E(\alpha)$.
3.5. Fourier decay of $\mu$
We now show that $\mu$ satisfies (1.2) if $\beta\geq 0$ and (1.5) if $\beta<0$.
Proposition 3.4.
If $10\leq\beta<$ 1, the measure $\mu$ satisfies
$|\widehat{\mu}(s)|\lesssim_{\epsilon}\log^{3}(|s|)|s|^{-\frac{1-\beta}{2+\alpha+\epsilon}},\qquad\text{for every $s\in\mathbb{Z}$}.$ (3.13)
If $-1<\beta<0$, the measure $\mu$ satisfies
$|\widehat{\mu}(s)|\lesssim_{\epsilon}\log^{3}(|s|)|s|^{-\frac{1}{2+\alpha+\epsilon}},\qquad\text{for every $s\in\mathbb{Z}$}.$ (3.14)
Proof.
First, we claim that for all $k\in\mathbb{N}$,
$|\widehat{\mu}_{k}(s)|\leq\log^{3}(|s|)(|s|^{-\frac{1}{2+\alpha+\epsilon_{k}}}+|s|^{-\frac{1-\beta}{2+\alpha+\epsilon_{k}}})\qquad\text{if}\ |s|\geq\min(M_{k},M_{k}^{1+\beta})/2.$ (3.15)
If $\min(M_{k},M_{k}^{1+\beta})/2\leq|s|\leq 2M_{k}^{2+\alpha+\epsilon_{k}}$, it follows from (3.9) that
$\begin{split}|\widehat{\mu}_{k}(s)|&\leq\log^{3}(|s|)M_{k}^{-1}(1+M_{k}^{\beta})\\ &\lesssim\log^{3}(|s|)(|s|^{-\frac{1}{2+\alpha+\epsilon_{k}}}+|s|^{-\frac{1-\beta}{2+\alpha+\epsilon_{k}}}).\end{split}$
If $|s|\geq 2M_{k}^{2+\alpha+\epsilon_{k}}$, (3.3) and (3.10) imply that
$\begin{split}|\widehat{\mu_{k}}(s)|&\leq\log(M_{k})\log(|s|)M_{k}^{-1+(2+\alpha)(N_{k}+1)}(1+M_{k}^{\beta})|s|^{-N_{k}}\\ &\leq\log^{2}(|s|)|s|^{\frac{-1+(2+\alpha)-\epsilon_{k}N_{k}}{2+\alpha+\epsilon_{k}}}(1+|s|^{\frac{\beta}{2+\alpha+\epsilon_{k}}}).\end{split}$
Since $\epsilon_{k}N_{k}\geq 2+\alpha$, we obtain
$|\widehat{\mu_{k}}(s)|\leq\log^{2}(|s|)(|s|^{-\frac{1}{2+\alpha+\epsilon_{k}}}+|s|^{-\frac{1-\beta}{2+\alpha+\epsilon_{k}}}).$
We have established (3.15).
For a given $\epsilon>0$, we choose $k$ such that $\epsilon_{k}\leq\epsilon$. For $\ell>k$, (3.15) implies that
$|\widehat{\mu}_{l}(s)|\lesssim\log^{3}(|s|)(|s|^{-\frac{1}{2+\alpha+\epsilon}}+|s|^{-\frac{1-\beta}{2+\alpha+\epsilon}})\qquad\text{if}\ |s|\geq\min(M_{l},M_{l}^{1+\beta})/2.$ (3.16)
If $\min(M_{i-1},M_{i-1}^{1+\beta})/2\leq|s|\leq\min(M_{i},M_{i}^{1+\beta})/2$ for some $k+1\leq i\leq l$, we use (3.8) repeatedly and obtain that
$\begin{split}|\widehat{\mu}_{l}(s)|&\leq|\widehat{\mu}_{i-1}(s)|+\sum_{j=i}^{l}\log(M_{j})M_{j}^{-96}\\ &\leq\log^{2}(|s|)(|s|^{-\frac{1}{2+\alpha+\epsilon}}+|s|^{-\frac{1-\beta}{2+\alpha+\epsilon}})+|s|^{-1}.\end{split}$ (3.17)
Combining (3.16) and (3.17), we get
$|\widehat{\mu}_{\ell}(s)|\lesssim\log^{3}(|s|)(|s|^{-\frac{1}{2+\alpha+\epsilon}}+|s|^{-\frac{1-\beta}{2+\alpha+\epsilon}})\qquad\text{if}\ |s|\geq\min(M_{k},M_{k}^{1+\beta})/2$
for any $\ell\geq k$. Letting $\ell\rightarrow\infty$, we obtain (3.13) and (3.14). Note that the implicit constant in (3.13) and (3.14) depends on $\epsilon$, since our choice of $k$ depends on $\epsilon$. ∎
Since $\operatorname*{supp}(\mu)\subseteq\mathbb{T}$, $\widehat{\mu}(\xi)=\mathcal{F}(\mu)(\xi)$ for integer $\xi$. So the preceding proposition shows that $\mu$ satisfies (1.2) and (1.5) when $\xi\in\mathbb{Z}$. In fact, by a standard argument (see, e.g., [20, pp.252-253] or [34, p.69]) the same estimates extend to all $\xi\in\mathbb{R}$. Thus we have proved $\mu$ satisfies (1.2) if $\beta\geq 0$ and (1.5) if $\beta<0$.
3.6. Regularity of $\mu$
We now show that $\mu$ satisfies (1.1) if $\beta\geq 0$ and (1.1) if $\beta\leq 0$. We start with some definitions and lemmas.
For $l>k$, define
$\mu_{l,k}:=g_{l}\cdots g_{k+1}\mu_{k-1}=g_{l}\cdots g_{k+1}g_{k-1}\cdots g_{1}\mu_{0}$
and $\mu_{k,k}:=\mu_{k-1}$. In other words, $\mu_{l,k}$ is defined by omitting $g_{k}$ from the product that defines $\mu_{l}$. By re-indexing, the proof of equation (3.7) and (3.8) in Lemma 3.3 yields the following lemma, which will be used in the proof of regularity of $\mu$ and the failure of the extension estimate (R).
Lemma 3.5.
$\left|\widehat{\mu}_{l,k}(s)\right|\leq 2\qquad\text{for every $s\in\mathbb{Z}$}$ (3.18)
and
$\left|\widehat{\mu}_{l,k}(s)-\widehat{\mu}_{l-1,k}(s)\right|\leq\log(M_{l})M_{l}^{-96}\qquad\text{if}\ |s|\leq\min(M_{l},M_{l}^{1+\beta})/2.$ (3.19)
To prove the regularity of the measure $\mu$, we define $k$-interval.
Definition 3.6.
If $k$ is odd, a $k$-interval is a connected component of $E(\alpha,k)\cup C(\beta,k)$. If $k$ is even, a $k$-interval is a connected component of $E(\alpha,k)$.
Note (i) the intervals of $E(\alpha,k)$ and of $C(\beta,k)$ are each of length $M_{k}^{-(2+\alpha)}$, (ii) the intervals of $E(\alpha,k)$ which do not contain the origin are $M_{k}^{-2}$-separated, and (iii) the intervals of $C(\beta,k)$ are $M_{k}^{-(1+\beta)}$-separated. Therefore each $k$-interval which does not contain the origin is a union of at most one interval of $E(\alpha,k)$ and at most one interval of $C(\beta,k)$. Also, observe that if $\beta=0$, each interval in $C(\beta,k)$ is in $E(\alpha,k)$, so a $k$-interval is just an interval of $E(\alpha,k)$ in this special case.
Lemma 3.7.
Recall that $F$ is a nonnegative Schwartz function supported on $\mathbb{T}$. Let $I$ be a $k$-interval centered at $x_{I}$ such that $|x_{I}|\geq 10M_{k}^{-(2+\alpha)}$ and define
$F_{I}(x)=F\left(\frac{1}{2|I|}(x-x_{I})\right).$
Then, for any $k\in\mathbb{N}$,
$|\widehat{g_{k}F_{I}}(s)|\lesssim_{N}\log(M_{k})M_{k}^{-2}(1+M_{k}^{-\beta})\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-N}.$
Proof.
For a $k$-interval $I$, let $2I$ be an interval of length $2|I|$ with the same center as $I$. The function $F_{I}$ is supported on an interval of length $2|I|\sim M_{k}^{-(2+\alpha)}$ and its support does not intersect a $k$-interval which contains the origin. Therefore, we have
$\sum_{p\in P_{M_{k}}}\sum_{|r|\leq(p-1)/2}F(M_{k}^{1+\alpha}(px-r))F_{I}(x)=\sum_{(p,r)\in P_{M_{k}}(I)}F(M_{k}^{1+\alpha}(px-r))F_{I}(x)$
for some set $P_{M_{k}}(I)$ which contains $\lesssim 1$ pairs $(p,r)$. Likewise, for odd $k$, we have
$\sum_{|a|\leq(B_{k}-1)/2}F(M_{k}^{\alpha-\beta+1}(B_{k}x-a))F_{I}(x)=\sum_{a\in B_{k}(I)}F(M_{k}^{\alpha-\beta+1}(B_{k}x-a))F_{I}(x)$
for some set $B_{k}(I)$ of size $\lesssim 1$. Therefore, we have for even $k$ that
$g_{k}F_{I}(x)=\frac{M_{k}^{1+\alpha}}{|P_{M_{k}}|}\sum_{(p,r)\in P_{M_{k}}(I)}F(M_{k}^{1+\alpha}(px-r))F_{I}(x),$
and we have for odd $k$ that
$\begin{split}g_{k}F_{I}(x)=&\frac{M_{k}^{1+\alpha}}{|P_{M}|+M_{k}^{\beta}+1}\sum_{(p,r)\in P_{M_{k}}(I)}F(M_{k}^{1+\alpha}(px-r))F_{I}(x)\\ &+\frac{M_{k}^{1+\alpha}(1+M_{k}^{-\beta})}{|P_{M}|+M_{k}^{\beta}+1}\sum_{a\in B_{k}(I)}F(M_{k}^{\alpha-\beta+1}(B_{k}x-a))F_{I}(x).\end{split}$
The functions
$F(M_{k}^{1+\alpha}(px-r))F_{I}(x)\qquad\text{and}\qquad F(M_{k}^{\alpha-\beta+1}(B_{k}x-a))F_{I}(x)$
are Schwartz functions whose supports are contained in subintervals of $\mathbb{T}$ of length $\sim M_{k}^{-(2+\alpha)}$. Thus, their Fourier coefficients at $s$ are bounded above by a constant multiple of
$M_{k}^{-(2+\alpha)}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-N}$
where the implicit constant depends on $N$. We then obtain for even $k$ that
$|\widehat{g_{k}F_{I}}(s)|\lesssim_{N}\frac{\log M_{k}}{M_{k}^{2}}\sum_{(p,r)\in P_{M_{k}}(I)}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-N},$
and for odd $k$ that
$|\widehat{g_{k}F_{I}}(s)|\lesssim_{N}\frac{\log(M_{k})}{M_{k}^{2}}\left(\sum_{(p,r)\in P_{M_{k}}(I)}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-N}+(1+M_{k}^{-\beta})\sum_{a\in B_{k}(I)}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-N}\right).$
Since the sizes of $P_{M_{k}}(I)$ and $B_{k}(I)$ are $\lesssim 1$, we establish the inequality. ∎
We will consider two kinds of $k$-intervals. Let $\mathcal{J}_{2}$ be the set of $k$-intervals $I$ such that $2I$ intersects $C(\beta,k)$ and let $\mathcal{J}_{1}$ be the remaining $k$-intervals. Of course, any $k$-interval in $\mathcal{J}_{1}$ will be an interval from $E(\alpha,k)$. Notice that if $k$ is even, then $\mathcal{J}_{2}$ is empty.
Lemma 3.8.
If $I$ is a $k$-interval in $\mathcal{J}_{1}$,
$\mu(I)\lesssim\log^{3}(M_{k})M_{k}^{-2},$ (3.20)
and
$\mu(I)\lesssim\log^{3}(M_{k-1})M_{k-1}^{\alpha}\log(M_{k})M_{k}^{-2}.$ (3.21)
If $I$ is a $k$-interval in $\mathcal{J}_{2}$,
$\mu(I)\lesssim\log^{3}(M_{k})M_{k}^{-2}(2+M_{k}^{-\beta}).$ (3.22)
Proof.
We consider $\mu_{k}$ first.
Estimate of $\mu_{k}$ for $k$-intervals in $\mathcal{J}_{1}$
Let $I\in\mathcal{J}_{1}$. We will find an upper bound on $\mu_{k}(2I)$. Observe that $\mu_{k}(2I)$ is given by the integral
[…]
The definition of $g_{k}$ and the fact that $F_{0}$ is supported away from the origin guarantee that on the support of $F_{0}$ for even values of $k$ we have the estimate
$g_{k}(x)\leq\frac{C_{0}M_{k}^{1+\alpha}}{|P_{M_{k}}|}.$
For odd values of $k$, $2I$ does not intersect $C(\beta,k)$ since $I\in\mathcal{J}_{1}$. Thus, the definition of $g_{k}$ and the support of $F_{0}$ guarantee that we have the estimate
$g_{k}(x)\leq\frac{C_{0}M_{k}^{1+\alpha}}{|P_{M_{k}}|+M_{k}^{\beta}+1}\qquad\text{for}\ x\in 2I.$
In either case, we have the estimate
$g_{k}(x)\leq C_{0}D\log(M_{k})M_{k}^{\alpha}\qquad\text{for}\ x\in 2I.$ (3.23)
From (3.3), we get $C_{0}D\leq\log(M_{k})$. Therefore,
$g_{k}(x)\leq\log^{2}(M_{k})M_{k}^{\alpha}\leq M_{k}^{1+\alpha}.$
By (3.2), $\log(M_{k})M_{k}^{1+\alpha}\leq\log(M_{k+1})$. For $x\in 2I$, we get
$\begin{split}g_{1}\cdots g_{k}(x)&\leq(M_{1}M_{2}\cdots M_{k-1})^{1+\alpha}\log^{2}(M_{k})M_{k}^{\alpha}\\ &\leq\log(M_{2})M_{2}^{1+\alpha}(M_{3}\cdots M_{k-1})^{1+\alpha}\log^{2}(M_{k})M_{k}^{\alpha}\\ &\leq\log(M_{3})M_{3}^{1+\alpha}(M_{4}\cdots M_{k-1})^{1+\alpha}\log^{2}(M_{k})M_{k}^{\alpha}\\ &\qquad\vdots\\ &\leq\log^{3}(M_{k})M_{k}^{\alpha}.\end{split}$
Thus, we obtain
[…]
Since $k$-intervals have Lebesgue measure $\sim M_{k}^{-(2+\alpha)}$, we have established the bound
$\mu_{k}(2I)\lesssim\log^{3}(M_{k})M_{k}^{-2}.$ (3.24)
Also, by (3.23) we have that
[…]
Estimate of $\mu_{k}$ for $k$-intervals in $\mathcal{J}_{2}$
Now, let $I\in\mathcal{J}_{2}$. We will find an upper bound on $\mu_{k}(2I)$. Since $\mathcal{J}_{2}$ is nonempty, $k$ must be odd. Instead of (3.23), we have
$g_{k}(x)\leq C_{0}D\log(M_{k})M_{k}^{\alpha}(2+M_{k}^{-\beta})\qquad\text{for}\ x\in 2I.$
Similarly, we get
$g_{1}\cdots g_{k}(x)\leq\log^{3}(M_{k})M_{k}^{\alpha}(2+M_{k}^{-\beta})\qquad\text{for}\ x\in 2I.$
Therefore,
$\mu_{k}(2I)\lesssim\log^{3}(M_{k})M_{k}^{-2}(2+M_{k}^{-\beta}).$ (3.26)
Estimate of $\mu$ for $k$-intervals
Let $I$ be a $k$-interval. We seek to estimate $\mu_{l}(I)$ for $l>k$. Let $F_{I}$ be the nonnegative Schwartz function defined in Lemma 3.7 and write
[…]
We apply Plancherel’s theorem to write
$\mu_{l}(F_{I})=\sum_{s\in\mathbb{Z}}\widehat{g_{k}F_{I}}(s)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(s)\qquad\text{and}\qquad\mu_{k}(F_{I})=\sum_{s\in\mathbb{Z}}\widehat{g_{k}F_{I}}(s)\widehat{\mu}_{k-1}(s).$
Let us consider
$\mu_{l}(F_{I})-\mu_{k}(F_{I})=\sum_{s\in\mathbb{Z}}\widehat{g_{k}F_{I}}(s)((g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(s)-\widehat{\mu}_{k-1}(s)).$
If $|s|\geq M_{k}^{10(2+\alpha)}$, we use the the decay of $\widehat{g_{k}F_{I}}$. By (3.7) and (3.18) we have
$|(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(s)-\widehat{\mu}_{k-1}(s)|\lesssim 1.$
By Lemma 3.7 with $N=10$, we obtain that
$\begin{split}\sum_{|s|\geq M_{k}^{10(2+\alpha)}}&|\widehat{g_{k}F_{I}}(s)((g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(s)-\widehat{\mu}_{k-1}(s))|\\ &\lesssim\log(M_{k})M_{k}^{-2}(1+M_{k}^{-\beta})\sum_{|s|\geq M_{k}^{10(2+\alpha)}}\left(\frac{M_{k}^{2+\alpha}}{|s|}\right)^{10}\lesssim M_{k}^{-80(2+\alpha)}.\end{split}$
If $|s|\leq M_{k}^{10(2+\alpha)}$, we use the estimate for $(g_{l}\cdots g_{k+1}g_{k-1}\cdots g_{1}F_{0})^{\wedge}(s)$. Since $M_{k}^{10(2+\alpha)}\leq\min(M_{k+1},M_{k+1}^{1+\beta})/2$, we use (3.19) to obtain
$\begin{split}|(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(s)-\widehat{\mu}_{k-1}(s)|&\leq\sum_{j=k+1}^{l}\log(M_{j})M_{j}^{-96}\\ &\lesssim M_{k+1}^{-90}.\end{split}$
By Lemma 3.7 with $N=0$,
$\begin{split}\sum_{|s|\leq M_{k}^{10(2+\alpha)}}&|\widehat{g_{k}F_{I}}(s)((g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(s)-\widehat{\mu_{k-1}}(s))|\\ &\lesssim\log(M_{k})M_{k}^{-2+10(2+\alpha)}(1+M_{k}^{-\beta})M_{k+1}^{-90}\leq M_{k}^{-80(2+\alpha)}.\end{split}$
Combining the two cases, we have $\mu_{l}(F_{I})\lesssim\mu_{k}(F_{I})+M_{k}^{-80(2+\alpha)}$. Since $F_{I}$ is supported on $2I$, we have $\mu_{k}(F_{I})\lesssim\mu_{k}(2I)$. Thus, we get
$\mu_{l}(F_{I})\lesssim\mu_{k}(2I)+M_{k}^{-80(2+\alpha)}.$ (3.27)
Because $F_{I}$ is a continuous function and $\mu_{l}$ converges weakly to $\mu$, we conclude that $\mu(F_{I})\lesssim\mu_{k}(2I)+M_{k}^{-80(2+\alpha)}$. Furthermore, since $1\lesssim F_{I}$ on $I$, we conclude that $\mu(I)\lesssim\mu(F_{I})$. So we can conclude
$\mu(I)\lesssim\mu_{k}(2I)+M_{k}^{-80(2+\alpha)}.$ (3.28)
Combining (3.28) with (3.24), (3.25) and (3.26) implies the estimates (3.20), (3.21) and (3.22) respectively.
∎ Now we are ready to prove the regularity of the measure $\mu$. Recall that $\mu$ is the weak limit of a subsequence of the measures $\mu_{k}$.
Proposition 3.9.
For any interval $I\subseteq\mathbb{T}$, if $10\leq\beta<$ 1,
$\mu(I)\lesssim\log^{3}(|I|^{-1})|I|^{\frac{2}{2+\alpha}}.$
If $-1<\beta<0$,
$\mu(I)\lesssim\log^{3}(|I|^{-1})|I|^{\frac{2+\beta}{2+\alpha}}.$
Proof.
Assume that $I$ is an arbitrary interval satisfying
$M_{k}^{-(2+\alpha)}\leq|I|\leq M_{k-1}^{-(2+\alpha)}.$
We will estimate $\mu(I)$ by counting the number of $k$-intervals that intersect $I$. We have the decomposition
$\mu(I)\leq\sum_{J\in\mathcal{J}_{1}}\mu(J)+\sum_{J\in\mathcal{J}_{2}}\mu(J).$
Observe that the sum over $J\in\mathcal{J}_{2}$ is empty if $k$ is even.
We will first estimate the sum over those $J\in\mathcal{J}_{1}$. The number of $k$-intervals $J\in\mathcal{J}_{1}$ intersecting $I$ can be estimated in a manner similar to that of Papadimitropoulos [26]. We will count the rational numbers $\frac{r}{p}$ contained in the interval $3I$ where $p\in P_{M_{j}}$. Because each $k$-interval intersecting $I$ must have its center contained in $3I$, this will give an upper bound on the number of $k$-intervals intersecting $I$.
We will split into two cases based on whether the interval $I$ is larger than $M_{k}^{-1}$.
Case 1: $M_{k}^{-(2+\alpha)}\leq|I|\leq M_{k}^{-1}$
Since the rational numbers in $\{r/p:p\in P_{M_{k}}\}$ are easily seen to be $M_{k}^{-2}$-separated, the number of rational numbers of this form contained in $3I$ is bounded above by $\max(3|I|M_{k}^{2},1)$. Since each such $k$ interval has $\mu$-measure at most a constant times $\log^{3}(M_{k})M_{k}^{-2}$ by (3.20), we obtain the following bound:
$\sum_{J\in\mathcal{J}_{1}}\mu(J)\lesssim\log^{3}(M_{k})\max(|I|,M_{k}^{-2}).$ (3.29)
We will now show that this implies the desired regularity estimate for $\sum_{J\in\mathcal{J}_{1}}\mu(J)$. If the maximum on the right side of (3.29) is realized by $\log^{3}(M_{k})|I|$, then we use the inequality $|I|\leq M_{k}^{-1}$ and conclude
$\sum_{J\in\mathcal{J}_{1}}\mu(J)\lesssim\log^{3}(|I|^{-1})|I|\leq\log^{3}(|I|^{-1})|I|^{\frac{2}{2+\alpha}}.$
If the maximum in (3.29) is realized by $\log^{3}(M_{k})M_{k}^{-2}$, we use the inequality $M_{k}^{-(2+\alpha)}\leq|I|$ to conclude
$\sum_{J\in\mathcal{J}_{1}}\mu(J)\lesssim\log^{3}(|I|^{-1})|I|^{\frac{2}{2+\alpha}}.$
This completes the proof of the regularity estimate for $\sum_{J\in\mathcal{J}_{1}}\mu(J)$ when $|I|$ is small.
Case 2: $M_{k}^{-1}\leq|I|\leq M_{k-1}^{-(2+\alpha)}$
We will count the number of rational numbers of the form $\{r/p:p\in P_{M_{k}}\}$ contained in $3I$ in a different way. For some fixed $p_{0}\in P_{M_{k}}$, the total number of rational numbers $r/p_{0}$ contained in $3I$ is at most $\lesssim M_{k}|I|$. We get that the total number of rational numbers $\{r/p:p\in P_{M_{k}}\}$ is bounded above by a constant times $|I||P_{M_{k}}|M_{k}$. Using (3.21) and $|P_{M_{k}}|\sim\frac{M_{k}}{\log M_{k}}$, we obtain the following bound:
$\begin{split}\sum_{J\in\mathcal{J}_{1}}\mu(J)&\lesssim\log^{3}(M_{k-1})M_{k-1}^{\alpha}|P_{M_{k}}|\log(M_{k})M_{k}^{-1}|I|\\ &\lesssim\log^{3}(M_{k-1})M_{k-1}^{\alpha}|I|.\end{split}$ (3.30)
Combining (3.30) with the inequality $|I|\leq M_{k-1}^{-(2+\alpha)}$, we conclude
$\sum_{J\in\mathcal{J}_{1}}\mu(J)\lesssim\log^{3}(|I|^{-1})|I|^{\frac{2}{2+\alpha}}.$
It remains to control the contribution of those $J\in\mathcal{J}_{2}$. Note that this step is only necessary if $k$ is odd. If a $k$-interval $J\in\mathcal{J}_{2}$ intersects $I$, it follows that the subinterval from $C(\beta,k)$ which intersects $2J$ will be contained in $6I$ and for each subinterval from $C(\beta,k)$, it can intersect at most $2$ $k$-intervals $J\in\mathcal{J}_{2}$. So, it is sufficient to count the number of intervals $C(\beta,k)$ in $6I$.
To count the number of intervals of $C(\beta,k)$ in $6I$, we use the $M^{-(1+\beta)}$-separation of these intervals. This separation implies that the number of such intervals in $6I$ is at most a constant times $\max(M_{k}^{1+\beta}|I|,1)$. Since the $\mu$-measure of each $k$-interval is bounded above by a constant times $\log^{3}(M_{k})M_{k}^{-2}(2+M_{k}^{-\beta})$ by (3.22), we obtain
$\begin{split}\sum_{J\in\mathcal{J}_{2}}\mu(J)&\lesssim\log^{3}(M_{k})M_{k}^{-2}(2+M_{k}^{-\beta})\max(M_{k}^{1+\beta}|I|,1).\end{split}$ (3.31)
If the maximum on the right side of (3.31) is realized by $M_{K}^{1+\beta}|I|$, by the inequality $M_{k}^{-(2+\alpha)}\leq|I|$, we get
$\begin{split}\sum_{J\in\mathcal{J}_{2}}\mu(J)&\lesssim\log^{3}(|I|^{-1})(|I|^{\frac{3+\alpha-\beta}{2+\alpha}}+|I|^{\frac{3+\alpha}{2+\alpha}})\\ &\lesssim\log^{3}(|I|^{-1})(|I|^{\frac{2}{2+\alpha}}+|I|^{\frac{2+\beta}{2+\alpha}}).\end{split}$
If the maximum on the right side of (3.31) is realized by $1$, by the inequality $M_{k}^{-(2+\alpha)}\leq|I|$, we conclude
$\sum_{J\in\mathcal{J}_{2}}\mu(J)\lesssim\log^{3}(|I|^{-1})(|I|^{\frac{2}{2+\alpha}}+|I|^{\frac{2+\beta}{2+\alpha}}).$
Hence $\mu(I)\lesssim\log^{3}(|I|^{-1})|I|^{\frac{2}{2+\alpha}}$ if $10\leq\beta<$ 1 and $\mu(I)\lesssim\log^{3}(|I|^{-1})|I|^{\frac{2+\beta}{2+\alpha}}$ if $-1<\beta<0$, as desired. ∎
4. Failure of the extension estimate
In this section, we define a sequence of function $f_{k}$ which witness the failure of the estimate (R). More specifically, we shall prove (1.3) if $\beta\geq 0$ and (1.6) if $\beta<0$. This will complete the proofs of Theorem 1.7 and Theorem 1.8.
4.1. Definitions and Lemmas
We define a sequence of functions $\{f_{k}\}_{\begin{subarray}{c}k\in\mathbb{N}\\ k\text{ odd}\end{subarray}}$ by
$f_{k}(x)=\sum_{|a|\leq(B_{k}-1)/2}F(M_{k}^{\alpha-\beta+1}(B_{k}x-a)).$
To obtain a lower bound of $\widehat{f_{k}\mu_{l}}(s)$ for $l>k$, we write
$\widehat{f_{k}\mu_{l}}(s)=\sum_{t\in\mathbb{Z}}\widehat{f_{k}g_{k}}(s-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)$
and split $f_{k}g_{k}$ into two functions. Let us consider
$h_{1}(x)=\frac{M_{k}^{1+\alpha}(1+M_{k}^{-\beta})}{|P_{M_{k}}|+M_{k}^{\beta}+1}\sum_{|a|\leq(B_{k}-1)/2}F^{2}(M_{k}^{\alpha-\beta+1}(B_{k}x-a))$
and
$h_{2}(x)=\frac{M_{k}^{1+\alpha}}{|P_{M_{k}}|+M_{k}^{\beta}+1}\sum_{p\in P_{M_{k}}}\sum_{|r|\leq(p-1)/2}F(M_{k}^{1+\alpha}(px-r))\sum_{|a|\leq(B_{k}-1)/2}F(M_{k}^{\alpha-\beta+1}(B_{k}x-a)).$
Then, $g_{k}f_{k}=h_{1}+h_{2}$ and we have
$\widehat{h}_{1}(s)=\frac{1+M_{k}^{\beta}}{|P_{M_{k}}|+M_{k}^{\beta}+1}\widehat{F^{2}}\left(\frac{s}{B_{k}M_{k}^{1+\alpha-\beta}}\right)\mathbf{1}_{B_{k}\mathbb{Z}}(s)$
and
$\widehat{h}_{2}(s)=\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|+M_{k}^{\beta}+1}\sum_{p\in P_{M_{k}}}\sum_{t\in\mathbb{Z}}\widehat{F}\left(\frac{t}{B_{k}M_{k}^{1+\alpha-\beta}}\right)\widehat{F}\left(\frac{s-t}{pM_{k}^{1+\alpha}}\right)\mathbf{1}_{B_{k}\mathbb{Z}}(t)\mathbf{1}_{p\mathbb{Z}}(s-t).$
We will use the following estimate on $\widehat{h}_{2}(s)$.
Lemma 4.1.
If $10<\beta<$ 1,
$|\widehat{h}_{2}(s)|\lesssim\frac{\log(M_{k})}{M_{k}}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.$ (4.1)
If $-1<\beta<0$,
$|\widehat{h}_{2}(s)|\lesssim\frac{1}{M_{k}}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.$ (4.2)
Proof.
First, let us consider when $10<\beta<$ 1. Note that there are at most $\frac{\log(|s-t|)}{\log M_{k}}$ values of $p\in P_{M_{k}}$ such that $p|(s-t)$. For fixed $s$, if $|s-t|\geq|s|/2$,
$\begin{split}\left|\sum_{p\in P_{M_{k}}}\widehat{F}\left(\frac{s-t}{pM_{k}^{1+\alpha}}\right)\mathbf{1}_{p\mathbb{Z}}(s-t)\right|&\lesssim\frac{\log(|s-t|)}{\log M_{k}}\left(1+\frac{|s-t|}{M_{k}^{2+\alpha}}\right)^{-11}\\ &\lesssim\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.\end{split}$
If $|s-t|\leq|s|/2$, then $|s|\sim|t|$. Thus,
$\left|\widehat{F}\left(\frac{t}{B_{k}M_{k}^{1+\alpha-\beta}}\right)\right|\lesssim\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\left(1+\frac{|t|}{M_{k}^{2+\alpha}}\right)^{-10}$
and
$\begin{split}\left|\sum_{p\in P_{M_{k}}}\widehat{F}\left(\frac{s-t}{pM_{k}^{1+\alpha}}\right)\mathbf{1}_{p\mathbb{Z}}(s-t)\right|&\lesssim\frac{\log(|s-t|)}{\log M_{k}}\left(1+\frac{|s-t|}{M_{k}^{2+\alpha}}\right)^{-1}\\ &\lesssim 1.\end{split}$
Therefore, we have
$|\widehat{h}_{2}(s)|\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{t\in\mathbb{Z}}\left(1+\frac{|t|}{M_{k}^{2+\alpha}}\right)^{-10}\mathbf{1}_{B_{k}\mathbb{Z}}(t).$
If $\mathbf{1}_{B_{k}\mathbb{Z}}(t)\neq 0$, then $t=B_{k}m$ for some $m\in\mathbb{Z}$. Thus,
$\begin{split}|\widehat{h}_{2}(s)|&\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{m\in\mathbb{Z}}\left(1+\frac{|B_{k}m|}{M_{k}^{2+\alpha}}\right)^{-10}\\ &\lesssim\frac{\log(M_{k})}{M_{k}}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.\end{split}$
Thus, we obtain (4.1)
Let us consider when $-1<\beta<0$. For fixed $s$, if $|s-t|\geq|s|/2$,
$\left|\widehat{F}\left(\frac{s-t}{B_{k}M_{k}^{1+\alpha-\beta}}\right)\right|\lesssim\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.$
If $|s-t|\leq|s|/2$, then $|s|\sim|t|$. Thus,
$\left|\widehat{F}\left(\frac{t}{B_{k}M_{k}^{1+\alpha-\beta}}\right)\right|\lesssim\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\left(1+\frac{|t|}{M_{k}^{2+\alpha}}\right)^{-10}.$
Therefore, we have
$|\widehat{h}_{2}(s)|\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{p\in P_{M_{k}}}\sum_{t\in\mathbb{Z}}\left(1+\frac{|t|}{M_{k}^{2+\alpha}}\right)^{-10}\mathbf{1}_{B_{k}\mathbb{Z}}(t)\mathbf{1}_{p\mathbb{Z}}(s-t).$
Assume that $\mathbf{1}_{B_{k}\mathbb{Z}}(t)\mathbf{1}_{p\mathbb{Z}}(s-t)\neq 0$. Then $t$ satisfies the congruences
[…]
$t$ $\equiv s\quad\text{(mod $p$)}.$
Moreover, we have assumed that $B_{k}$ is relatively prime to $p$. Therefore, the Chinese remainder theorem implies that $t$ satisfies a congruence of the form
$t\equiv a_{s,p}B_{k}\text{(mod $B_{k}p$)}$
for some $0\leq a_{s,p}<p$. Thus $t=B_{k}pm+B_{k}a_{s,p}$ for some $m\in\mathbb{Z}$. Thus,
$\begin{split}|\widehat{h}_{2}(s)|&\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{p\in P_{M_{k}}}\sum_{m\in\mathbb{Z}}\left(1+\frac{|pB_{k}m+B_{k}a_{s,p}|}{M_{k}^{2+\alpha}}\right)^{-10}\\ &\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{p\in P_{M_{k}}}\sum_{m\in\mathbb{Z}}\left(1+\frac{|m+a_{s,p}/p|}{M_{k}^{\alpha-\beta}}\right)^{-10}\\ &\lesssim\frac{1}{M_{k}}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.\end{split}$
In the last inequality, we used that $\beta<0<\alpha$. Therefore, we obtain (4.2).
Lemma 4.2.
Let $l>k$. For $|m|\leq M_{k}^{1+\alpha-\beta}/10$, we have
$|\widehat{f_{k}\mu_{l}}(B_{k}m)|\gtrsim\log(M_{k})M_{k}^{-1}(1+M_{k}^{\beta}).$ (4.3)
Proof.
We claim that the following estimates hold for sufficiently large $k$ and for $|m|\leq M_{k}^{1+\alpha-\beta}/10$:
$|\widehat{h}_{1}(B_{k}m)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(0)|\gtrsim\log(M_{k})M_{k}^{-1}(1+M_{k}^{\beta});$ (4.4)
$\sum_{t\neq 0}|\widehat{h}_{1}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{3}(M_{k})M_{k}^{-99}(1+M_{k}^{\beta});$ (4.5)
$\sum_{t\in\mathbb{Z}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{2}(M_{k})M_{k}^{-1}\qquad\text{if}\ 0<\beta<1;$ (4.6)
$\sum_{t\in\mathbb{Z}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{1/10}(M_{k})M_{k}^{-1}\qquad\text{if}\ -1<\beta<0.$ (4.7)
By combining (4.4), (4.5), (4.6) and (4.7) we get (4.3) when $\beta\neq 0$.
If $\beta=0$, it suffices to consider $h_{1}$ only. Since $\beta=0$, let $B_{k}=p\in P_{M_{k}}$. We obtain
$\begin{split}h_{2}=&\frac{M_{k}^{1+\alpha}}{|P_{M_{k}}|+2}\sum_{|r|\leq(p-1)/2}F^{2}(M_{k}^{1+\alpha}(px-r))\\ &+\frac{M_{k}^{1+\alpha}}{|P_{M_{k}}|+2}\sum_{\begin{subarray}{c}p^{\prime}\in P_{M_{k}}\\ p^{\prime}\neq p\end{subarray}}\sum_{|r|\leq(p^{\prime}-1)/2}F(M_{k}^{1+\alpha}(p^{\prime}x-r))\sum_{|a|\leq(p-1)/2}F(M_{k}^{1+\alpha}(px-a)).\end{split}$
Let $h_{2,1}$ and $h_{2,2}$ denote the first term and the second term respectively. Then, $h_{2,1}$ is a bounded multiple of $h_{1}$. The function $h_{2,2}$ is nonzero only when $|x|\lesssim M_{k}^{-(2+\alpha)}$ in $\mathbb{T}$ since rational numbers in $\{r/p^{\prime},a/p:p^{\prime}\in P_{M_{k}},p\neq p^{\prime}\}$ are $M_{k}^{-2}$-separated. However, $\mu_{k-1}$ is supported away from the origin since $F_{0}=0$ if $|x|\leq 1/8$. Therefore, $h_{2,2}\cdot(g_{l}\cdots g_{k+1}\mu_{k-1})=0$. By (4.4) and (4.5), we obtain (4.3) when $\beta=0$.
Proof of (4.4)
Since $F$ is a Schwartz function such that $F(x)\gtrsim 1$ if $|x|\leq 1/4$, we have
$|\widehat{h}_{1}(B_{k}m)|\gtrsim\log(M_{k})M_{k}^{-1}(1+M_{k}^{\beta}).$
By iterating (3.8) and (3.19), we get
$|(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(0)-\widehat{F}_{0}(0)|\leq\sum_{i=1}^{l}\log(M_{i})M_{i}^{-96}\leq M_{1}^{-90}.$ (4.8)
Since $\widehat{F}_{0}(0)=1$, we have $|(g_{l}\cdots g_{k+1}g_{k-1}\cdots g_{1}F_{0})^{\wedge}(0)|\gtrsim 1$. Thus, (4.4) follows.
Proof of (4.5)
The proof is similar to that of the estimate on $\mu$ for $k$-intervals, but we also use the properties of $\widehat{h}_{1}$.
If $|t|\geq M_{k}^{10(2+\alpha)}$, $|B_{k}m-t|\geq|t|/2$ holds since $|B_{k}m|\lesssim M_{k}^{2+\alpha}$. Thus, we get
$|\widehat{h}_{1}(B_{k}m-t)|\lesssim\log(M_{k})M_{k}^{-1}(1+M_{k}^{\beta})\left(\frac{M_{k}^{2+\alpha}}{|t|}\right)^{10}.$
By (3.18), we obtain
$\sum_{|t|\geq M_{k}^{10(2+\alpha)}}|\widehat{h}_{1}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log(M_{k})(1+M_{k}^{\beta})M_{k}^{-1-80(2+\alpha)}.$
If $|t|\leq M_{k}^{10(2+\alpha)}$, by iterating (3.19), we have
$|(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim|\widehat{\mu}_{k-1}(t)|+M_{k+1}^{-90}.$
Also, observe that $t=B_{k}m$ for some nonzero $m$ since $\widehat{h}_{1}$ is supported on $B_{k}\mathbb{Z}$ and $t\neq 0$. Thus, $|t|\geq M_{k}^{1+\beta}\geq 2M_{k-1}^{2+\alpha}$, and (3.10) with $N=N(\beta)$ together with (3.2) imply that
$\begin{split}|\widehat{\mu}_{k-1}(t)|&\leq\log(M_{k-1})M_{k-1}^{-1+(2+\alpha)(N(\beta)+1)}(1+M_{k-1}^{\beta})\log(|t|)|t|^{-N(\beta)}\\ &\lesssim\log(M_{k})\log(|t|)|t|^{-N(\beta)}.\end{split}$
Using $|\widehat{h}_{1}(s)|\lesssim\log(M_{k})M_{k}^{-1}(1+M_{k}^{\beta})$, we obtain
$\begin{split}\sum_{0<|t|\leq M_{k}^{10(2+\alpha)}}&|\widehat{h}_{1}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\\ &\lesssim\log(M_{k})M_{k}^{-1}(1+M_{k}^{\beta})\left(M_{k}^{10(2+\alpha)}M_{k+1}^{-90}+\log(M_{k})\sum_{|t|\geq M_{k}^{1+\beta}}\log(|t|)|t|^{-N(\beta)}\right)\\ &\lesssim\log^{3}(M_{k})M_{k}^{-99}(1+M_{k}^{\beta}).\end{split}$
In the first inequality, we used that $|t|\geq M_{k}^{1+\beta}$ because of the support of $\widehat{h}_{1}$. Hence, we get (4.5).
Proof of (4.6) and (4.7)
Let us consider when $10<\beta<$ 1. By the same method we used in the proof of (4.5), we use Lemma 4.1 and obtain
$\sum_{|t|\geq M_{k}^{10(2+\alpha)}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log(M_{k})M_{k}^{-1-80(2+\alpha)}.$
By (3.2), we have $\log(M_{k})^{1/10}\geq 2M_{k-1}^{2+\alpha}$. By Lemma 4.1, $|\widehat{h}_{2}(s)|\lesssim\log(M_{k})M_{k}^{-1}$. Thus, we use (3.10) with $N=99$ and (3.19) to obtain
$\begin{split}\sum_{\log^{1/10}(M_{k})\leq|t|\leq M_{k}^{10(2+\alpha)}}&|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\\ &\lesssim\log(M_{k})M_{k}^{-1}\left(M_{k}^{10(2+\alpha)}M_{k+1}^{-90}+\log(M_{k})\sum_{|t|\geq\log^{1/10}(M_{k})}\log(|t|)|t|^{-99}\right)\\ &\lesssim\left(\log^{5}(M_{k})M_{k}\right)^{-1}.\end{split}$
By Lemma 4.1 and (3.18), we get
$\sum_{|t|\leq\log(M_{k})^{1/10}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{2}(M_{k})M_{k}^{-1}.$ (4.9)
Combining the three cases, we get (4.6).
When $-1<\beta<0$, we the proof is same except that $\log(M_{k})M_{k}^{-1}$ is replaced by $M_{k}^{-1}$ because of Lemma 4.1. Also, instead of (4.9), we have
$\sum_{|t|\leq\log(M_{k})^{1/10}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{1/10}(M_{k})M_{k}^{-1}.$
Therefore, we obtain (4.7).
4.2. Proof of (1.3) and (1.6)
Now we complete the proofs of Theorem 1.7 and Theorem 1.8 by proving (1.3) for $\beta\geq 0$ and (1.6) for $\beta<0$.
Let us prove (1.3) first. Assume $\beta\geq 0$. Observe that each $F(M_{k}^{\alpha-\beta+1}(B_{k}x-a))$ is supported on a $k$-interval in $\mathcal{J}_{2}$ and their supports are disjoint. Since $|F|\lesssim 1$ and $\beta\geq 0$, (3.22) implies that
${\left\|f_{k}\right\|}_{L^{q}(\mu)}^{q}\lesssim\log^{3}(M_{k})M_{k}^{-2}B_{k}\lesssim\log^{3}(M_{k})M_{k}^{\beta-1}\qquad\text{for}\ 1\leq q<\infty.$ (4.10)
We will estimate $\mathcal{F}(f_{k}\mu)(\xi+s)$ for $s\in\mathbb{Z}$ and $|\xi|\leq(\log(M_{k}))^{-3}$. In fact, since $\mu$ is supported on $[-1/2,1/2]$, we obtain
[…]
$\lesssim$ $|\xi|\log^{3}(M_{k})M_{k}^{\beta-1}.$
In the last inequality, we used (4.10). The bound on $\xi$ guarantees that this is much smaller than $\log(M_{k})M_{k}^{\beta-1}$. By Lemma 4.2, for large $k$, we have that $\mathcal{F}(f_{k}\mu)(B_{k}m+\xi)\gtrsim\log(M_{k})M_{k}^{\beta-1}$ whenever $|m|\leq M_{k}^{1+\alpha-\beta}/10$ Thus, we obtain
${\left\|\mathcal{F}(f_{k}\mu)\right\|}_{L^{p}(\lambda)}^{p}\geq\int_{|\xi|\leq(\log M_{k})^{-3}}\sum_{s\in\mathbb{Z}}|\mathcal{F}(f_{k}\mu)(s+\xi)|^{p}\gtrsim M_{k}^{\alpha-\beta+1}\log(M_{k})^{-3}(\log(M_{k})M_{k}^{\beta-1})^{p}.$
Therefore, we get
$\frac{{\left\|\mathcal{F}(f_{k}\mu)\right\|}_{L^{p}(\lambda)}}{{\left\|f_{k}\right\|}_{L^{q}(\mu)}}\gtrsim(\log(M_{k}))^{1-\frac{3}{p}-\frac{3}{q}}M_{k}^{(1-\beta)(\frac{1}{q}-1)+\frac{\alpha-\beta+1}{p}}.$
If $p<p_{+}(q)$, the right hand side diverges to $\infty$ as $k\rightarrow\infty$. We have established (1.3).
The proof of (1.6) is similar. Since $\beta<0$, instead of (4.10), we have
${\left\|f_{k}\right\|}_{L^{q}(\mu)}^{q}\lesssim\log^{3}(M_{k})M_{k}^{-2-\beta}B_{k}\lesssim\log^{3}(M_{k})M_{k}^{-1}\qquad\text{for}\ 1\leq q<\infty.$
We use Lemma 4.2 when $-1<\beta<0$ and we obtain
${\left\|\mathcal{F}(f_{k}\mu)\right\|}_{L^{p}(\lambda)}^{p}\gtrsim M_{k}^{\alpha-\beta+1}\log(M_{k})^{-3}(\log(M_{k})M_{k}^{-1})^{p}.$
Thus,
$\frac{{\left\|\mathcal{F}(f_{k}\mu)\right\|}_{L^{p}(\lambda)}}{{\left\|f_{k}\right\|}_{L^{q}(\mu)}}\gtrsim(\log(M_{k}))^{1-\frac{3}{p}-\frac{3}{q}}M_{k}^{(\frac{1}{q}-1)+\frac{\alpha-\beta+1}{p}}.$
If $p<p_{-}(q)$, the right hand side diverges to $\infty$ as $k\rightarrow\infty$. We have established (1.6).
Declarations
Funding
The authors did not receive support from any organization for the submitted work.
Competing Interests
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data Availability Statement
We do not analyse or generate any datasets, because our work is a mathematical proof.